OpenStudy (anonymous):

Organized Counting Problem: 1. Ten finalists are competing in a race at the Canada games: a) In how many different orders can the competitors finish the race? \(10\ \times\ 9\ \times\ 8 ......\times\ 1=3628800\) b) How many ways could the gold, silver, and bronze medal be awarded? \(10\ \times\ 9\ \times\ 8=720\) c) One of the finalists is a friend from your town. How many of the possible finishes would include them winning a medal? d) How many possible finishes would leave your friend out of the medal standings? I need help with letter C.

OpenStudy (anonymous):

she could be first, then \(9\times 8\) or second \(9\times 8\) or third \(9\times 8\) so i believe the answer is \[3\times 9\times 8\]

OpenStudy (anonymous):

but i suck at these so get a second opinion

OpenStudy (anonymous):

why 3 though? what i mean is, should it be \(9 \times 8 \times 7\)? but i'm not sure though i'm really confused

OpenStudy (anonymous):

last one you got \(9\times 8\times 7\) right?

OpenStudy (anonymous):

three represents the 3 positions your friend could end up in first, second or third

OpenStudy (anonymous):

if she comes in first, then second has 9 choices and third has 8 if she comes is second, then first has 9 choices and third has 8 if she comes in third, then first has 9 choices, second has 8

OpenStudy (anonymous):

so \[3\times 9\times 8\]

OpenStudy (anonymous):

aah okay i get it. thanks! ^_^

OpenStudy (anonymous):

yw